# DPBridge: Latent Diffusion Bridge for Dense Prediction

### v1

@@ 5.4 Ablation Studies
To evaluate the effectiveness of each design component and training setting, we conduct ablation studies using depth estimation tasks on the NYUv2 and KITTI datasets. The results are summarized in Tab.3 and Tab.4, with detailed discussions provided below.
5.4.1 Design Components
Finetuning VS Training From Scratch. To assess whether adapting pretrained diffusion backbones to our DPBridge improves perception performance, we compare the difference between finetuning and training from scratch paradigm. As shown in the first row of Tab.3, fine-tuning significantly outperforms training from scratch under identical experimental settings. This difference is due to two factors: first, foundation models trained on large, diverse datasets capture universal visual priors, aiding knowledge transfer to dense prediction tasks. Second, training from scratch requires more iterations for convergence. These observations suggest that building DPBridge on pretrained backbones is indeed advantageous.
Distribution Alignment Normalization. Next, we evaluate whether aligning the marginal distributions between diffusion bridge process and standard diffusion process is beneficial. As shown in the second and fourth rows of Tab.3, incorporating a distribution alignment normalization step before inputting data into the noise prediction network noticeably improves performance.
Image Consistency Loss. Regarding the auxiliary image consistency loss, the comparison between the third and fourth rows in Tab.3 shows that while effective, its impact is less significant than that of distribution alignment normalization. This is probably because it primarily influences regions where masked and valid areas overlap, which constitute only a relatively small proportion of the data. However, when integrated with other design components, the overall performance is significantly improved.
5.4.2 Bridge Process Hyperparameters
Parameterization and Backbone Selection. Apart from the noise prediction scheme used in our main experiments, we also consider how the choice of parameterization combined with the backbone selection influences performance. Results in Tab.4 reveal two key findings: (1) v-parameterization performs significantly worse compared with other two options, highlighting the need for further research on deriving objectives specific to the diffusion bridge process, and (2) ”noise prediction + SDv1.5” achieves the best results, indicating that the parameterization during fine-tuning should align with that of the backbone to maximize knowledge transfer efficiency.
Number of Sampling Steps. We evaluate performance with varying numbers of sampling steps during inference. As shown in the lower part of Tab.4, increasing sampling steps initially improves perception performance rapidly before leveling off. Notably, compared to the results in Tab.1, where all other diffusion-based methods are evaluated with 50 steps setting, DPBridge achieves comparable performance with fewer steps, demonstrating the efficiency of our proposed method.
@@ 6 Conclusion
In this paper, we present DPBridge, a latent diffusion bridge model to tackle general dense prediction tasks. We establish a direct mapping between input RGB images and their corresponding dense maps in the latent space based on fully-tractable diffusion bridge process. In this way, the prior distribution is no longer constrained to uninformative noise, but instead can be chosen as more structurally correlated ones. In addition, we introduce effective finetuning strategies to adapt large pretrained diffusion backbone to our bridge model construction, leveraging its rich prior knowledge to boost training speed and achieve remarkable perception performance. Future research directions include deriving v-prediction parameterization of diffusion bridge process, and extending to other dense prediction tasks.
Supplementary Material
In the supplementary material, we’ll provide contents that are not included in the main paper due to page limitations. We first fill in the derivation details of the loss function. Afterwards we will provide the additional quantitative comparisons with more baseline methods. Finally, we will present qualitative examples that apply DPBridge on more dense prediction tasks.
@@ Appendix A Derivation Details
In this section, we’ll provide derivation details of the maximum likelihood loss presented in Sec.4.5.1 . Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Eqn.(5) , whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$. The last two term of the posterior can be retrieved directly through the marginal distribution of diffusion bridge’s forward process, but further derivation is required for the first term. Specifically, as specified in Eqn.(3)
@@
To derive the transition probability $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, we substitute $\mathbf{z}_{0}$ in Eqn.(14) with derivation from Eqn.(15), obtain
@@
Putting all these terms back to Eqn.(5) , we’ll have:
@@
whose results are equivalent to Eqn.(6) and Eqn.(7).
@@
which is the same as in Eqn.(8) and Eqn.(9) . This concludes our derivation.
Appendix B Additional Quantitative Results
Next, we provide additional quantitative results on depth estimation (Tab.5) and surface normal prediction (Tab.6) tasks. Compared with Tab.1 and Tab.2 presented in the main paper, we conduct comparisons with more baseline methods, including feed-forward approaches and diffusion-based methods that is trained on more advanced Stable Diffusion (SD) v2.1 backbone. In depth estimation, DPBridge consistently matches state-of-the-art methods, while demonstrating robust generalization across datasets. For example, it achieves 6.9 AbsRel, 95.7 % $\delta_{1}$ on NYUv2, and 7.5 AbsRel, 94.7% $\delta_{1}$ on ScanNet. Similarly, for surface normal prediction, DPBridge also delivers competitive results, achieving mean error of 18.1 and $11.25^{\circ}$ accuracy of $54.2\%$ on NYUv2, as well as maintaining strong performance on other datasets like iBims-1 (mean: 24.3, $11.25^{\circ}:51.2\%$) and Sintel (mean: 41.2, $11.25^{\circ}:19.4\%$) and. These results highlight DPBridge’s versatility to generalize well across tasks and benchmarks, being comparable to both feed-forward and diffusion-based methods while offering a unified, efficient paradigm to handle dense prediction tasks.
Method Model Type NYUv2 KITTI ETH3D ScanNet DIODE
[…]
DiverseDepth [57] 11.7 87.5 19.0 70.4 22.8 69.4 10.9 88.2 37.6 63.1
MiDaS [43] 11.1 88.5 23.6 63.0 18.4 75.2 12.1 84.6 33.2 71.5
LeReS [58] 9.0 91.6 14.9 78.4 17.1 77.7 9.1 91.1 27.1 76.6
Omnidata [19] 7.4 94.5 14.9 83.5 16.6 77.8 7.5 93.2 33.9 74.2
HDN [60] 6.9 94.8 11.5 86.7 12.1 83.3 8.0 93.6 24.6 78.0
DPT [44] 9.8 90.3 10.0 90.1 7.8 94.6 8.2 93.8 18.2 75.8
Marigold* [32] 6.0 95.9 10.5 90.4 7.1 95.1 6.9 94.5 31.0 71.2
DepthFM* [24] 6.5 95.6 8.3 93.4 - - - - 22.5 80.0
Geowizard* [20] 5.7 96.2 14.4 82.0 7.5 94.3 6.1 95.8 31.4 77.1
E2E-FT* [21] 5.4 96.4 9.6 92.1 6.4 95.9 5.8 96.5 30.3 77.6
Marigold [32] 8.6 93.6 13.8 82.6 9.3 93.2 10.4 89.7 33.6 73.9
Geowizard [20] 8.5 93.7 16.2 75.6 9.1 93.1 9.0 93.0 32.0 72.2
E2E-FT [21] 8.2 94.0 15.4 77.9 10.6 89.8 9.2 92.3 32.9 73.7
ControlNet [61] 45.5 54.6 55.2 40.2 30.5 58.8 25.7 68.3 52.1 50.1
InstructPix2Pix [6] 35.2 60.0 47.7 51.2 26.1 66.7 21.4 75.3 44.8 58.0
I2SB [37] 24.9 73.4 37.7 57.9 40.6 52.0 21.2 66.1 40.2 59.1
BBDM [36] 27.9 65.5 47.2 45.1 46.4 48.8 22.7 67.0 34.6 64.7
DPBridge (Ours) 6.9 95.7 11.2 87.3 8.1 93.2 7.5 94.7 29.7 78.6
Table 5: Quantitative comparison of DPBridge with state-of-the-art affine-invariant depth estimation methods on zero-shot benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The superscript * indicates that this model is trained based on Stable Diffusion v2.1, otherwise is based on Stable Diffusion v1.5.
Method Model Type NYUv2 ScanNet iBims-1 Sintel
[…]
EESNU [3] 16.2 58.6 - - 20.0 58.5 42.1 11.5
Omnidata v1 [19] 23.1 45.8 22.9 47.4 19.0 62.1 41.5 11.4
Omnidata v2 [31] 17.2 55.5 16.2 60.2 18.2 63.9 40.5 14.7
DSINE [2] 16.4 59.6 16.2 61.0 17.1 67.4 34.9 21.5
Marigold* [32] 18.8 55.9 17.7 58.8 18.4 64.3 39.1 14.9
Geowizard* [20] 19.1 49.5 17.3 53.7 19.5 61.6 40.4 13.2
E2E-FT* [21] 16.5 60.4 14.7 66.1 16.1 69.7 33.5 22.3
Marigold [32] 22.3 40.0 23.3 37.6 28.0 38.5 52.1 8.1
Geowizard [20] 21.7 44.2 22.0 42.1 28.2 37.9 45.8 11.0
E2E-FT [21] 17.7 57.8 17.4 57.4 18.0 64.6 43.9 14.1
DPBridge 18.1 54.2 18.0 54.5 24.3 51.2 41.2 19.4
Table 6: Quantitative comparison of DPBridge with state-of-the-art surface normal prediction methods on zero-shot benchmarks. All metrics are presented in percentage terms. FF, DM, and BM correspond to feed-forward, diffusion models, and bridge models. The superscript * indicates that this model is trained based on Stable Diffusion v2.1, otherwise is based on Stable Diffusion v1.5.
@@ Appendix B Additional Quantitative Results
Furthermore, we also evaluate the robustness of DPBridge system by perturbing the input image with different types and different level intensities of noise. The evaluation is conducted under both depth estimation and surface normal prediction tasks. We use NYUv2 and KITTI to report the depth estimation results, as for surface normal prediction, we report the results on NYUv2 and iBims-1. Tab. 7 reveals that the model’s performance degrades as noise intensity increases, but the model performance remains acceptable under moderate-level-intensity noise. Among noise types, uniform noise shows the mildest effect on degradation, while salt & pepper and high-intensity gaussian noise cause the most significant performance decline. In addition, the noise intensity has more impact on surface normal prediction than depth estimation, showcasing the inherent characteristics between different tasks. These results emphasize the model’s robustness under moderate noise but still expose its vulnerability to extreme or specific types of noise.
@@
Appendix C Additional Qualitative Results
To further validate the effectiveness and generalization of DPBridge, we provide qualitative results for various dense prediction tasks, including semantic segmentation, optical flow prediction, and style transfer, by framing them as continuous-valued image-to-image translation problems. For semantic segmentation, we convert discrete labels into color maps and use our bridge model to predict colors in RGB space instead of categorical labels. For optical flow, the bridge process is defined between the flow image and the first image in the input RGB pair. To accommodate this, we adapt the UNet architecture to incorporate both RGB images as conditions during inference and convert the flow map into RGB space, predicting corresponding colors rather than uv coordinates. The following figures (Fig.5, Fig.6, Fig.7, Fig.8) illustrate that DPBridge achieves visually satisfactory results across different tasks. Note that only qualitative assessments are presented here; detailed implementations and quantitative evaluations are left as directions for future research.
Figure 5: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape [14] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 6: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ [35] dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 7: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel [7] dataset. The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
Figure 8: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics [55] dataset.

### v2

@@ 5.4 Ablation Studies
To evaluate the effectiveness of each design component and training setting, we conduct ablation studies using depth estimation tasks on the NYUv2 and KITTI datasets. The results are summarized in Table 3 and Figure 5, with detailed discussions provided below.
Finetuning VS Training From Scratch. To assess whether adapting pretrained diffusion backbones improves perception performance, we conduct comparison between finetuning and training from scratch paradigm. As shown in the first row of Table 3, fine-tuning significantly outperforms training from scratch under identical experimental settings. This difference is due to two factors: first, foundation models trained on large, diverse datasets. They capture universal visual priors, aiding knowledge transfer to dense prediction tasks. Second, training from scratch requires more iterations for convergence. In addition, from the last two rows of Table 1 and Table 2, we can see that finetuning on SD2.1 exceeds that on SD1.5, suggesting that the quality of prior knowledge will also influence the performance. These observations indicate that building DPBridge on superior pretrained backbones is indeed advantageous.
Distribution Alignment Normalization. Next, we evaluate whether aligning the marginal distributions between diffusion bridge process and standard diffusion process is beneficial. As shown in the second and fourth rows of Table 3, incorporating a distribution alignment normalization step before inputting data into the noise prediction network noticeably improves performance.
Image Consistency Loss. Regarding the auxiliary image consistency loss, the comparison between the third and fourth rows in Table 3 shows that while effective, its impact is less significant than that of distribution alignment normalization. This is probably because it primarily influences regions where masked and valid areas overlap, which constitute only a relatively small proportion of the data. However, when integrated with other design components, the overall performance is significantly improved.
Number of Sampling Steps. We evaluate performance with varying numbers of sampling steps during inference. As shown in Figure 5, increasing sampling steps initially improves perception performance rapidly and reach saturation after 10 steps. Notably, compared to the results in Table 1, where all other diffusion-based methods are evaluated with 50 steps setting, DPBridge achieves comparable performance with fewer steps, demonstrating the efficiency of our proposed method.
@@ 6 Conclusion
In this paper, we present DPBridge, a latent diffusion bridge model to tackle general dense prediction tasks through generative pipeline. We establish direct mapping between input images and corresponding signal maps in the latent space based on fully-tractable diffusion bridge process. In this way, the prior distribution is no longer constrained to uninformative noise, but instead can be chosen as more structurally correlated ones. In addition, we introduce effective finetuning strategies to adapt large pretrained diffusion backbone to our bridge model construction, leveraging its rich prior knowledge to boost training speed and achieve remarkable perception performance. Future research directions include deriving other parameterization of diffusion bridge process, and extending to more dense prediction tasks.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning and we wish to spark further research on image generation and perception. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Derivation Details
In this section, we’ll provide derivation details of the maximum likelihood loss presented in Section 4.4.1. Its core is the posterior distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0},\mathbf{z}_{T})$ as in Equation 5, whose calculation involves three terms: $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, $q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{T})$, and $q(\mathbf{z}_{t}|\mathbf{z}_{0},z\mathbf{z}_{T})$. The last two term of the posterior can be retrieved directly through the marginal distribution of diffusion bridge’s forward process, but further derivation is required for the first term. Specifically, as specified in Equation 3:
@@
To derive the transition probability $q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{T})$, we substitute $\mathbf{z}_{0}$ in Equation 14 with derivation from Equation 15, obtain
@@
Putting all these terms back to Equation 5, we’ll have:
@@
whose results are equivalent to Equation 6 and Equation 7.
@@
which is the same as in Equation 8 and Equation 9. This concludes our derivation.
Appendix B Experiments
B.1 Network Architecture
The basic structure of our network follows the original design of Stable Diffusion. To better utilize visual information ($\mathbf{z}_{T}$ in our case), we concatenating both intermediate samples and input image along the feature dimension and send in the noise prediction network as a whole. To address the dimension incompatibility with pretrained backbone, the input channels of the network are thus doubled to accommodate the concatenated input. In addition, as suggested by (Ke et al., 2024), we duplicate the weight tensor of these accommodated layers and divide the values by two to preserve the input magnitude to subsequent layers. As DPBridge is an image-conditioned generative framework, we employ null-text embedding for all cross-attention layers to eliminate the impact of text prompts.
B.2 Additional Quantitative Results
@@
In this section, we further evaluate the robustness of DPBridge system by perturbing the input image with different types and different level intensities of noise. In this experiment setting, we use Stable Diffusion 1.5 as pretrained backbone. The evaluation is conducted using under both depth estimation and surface normal prediction tasks. We use NYUv2 and KITTI to report the depth estimation results, as for surface normal prediction, we report the results on NYUv2 and iBims-1. Table 4 reveals that the model’s performance degrades as noise intensity increases, but the model performance remains acceptable under moderate-level-intensity noise. Among noise types, uniform noise shows the mildest effect on degradation, while salt & pepper and high-intensity gaussian noise cause the most significant performance decline. In addition, the noise intensity has more impact on surface normal prediction than depth estimation, showcasing the inherent characteristics between different tasks. These results emphasize the model’s robustness under moderate noise but still expose its vulnerability to extreme or specific types of noise.
@@
B.3 Additional Qualitative Results
To further validate the effectiveness and generalization of DPBridge, we provide qualitative results for various dense prediction tasks, including semantic segmentation, optical flow prediction, and style transfer, by framing them as continuous-valued image-to-image translation problems. For semantic segmentation, we convert discrete labels into color maps and use our bridge model to predict colors in RGB space instead of categorical labels. For optical flow, the bridge process is defined between the flow image and the first image in the input RGB pair. To accommodate this, we adapt the UNet architecture to incorporate both RGB images as conditions during inference and convert the flow map into RGB space, predicting corresponding colors rather than uv coordinates. The following figures (Figure 6, Figure 7, Figure 8, Figure 9) illustrate that DPBridge achieves visually satisfactory results across different tasks. Note that only qualitative assessments are presented here; detailed implementations and quantitative evaluations are left as directions for future research.
Figure 6: Qualitative examples of applying DPBridge to semantic segmentation task on CityScape (Cordts et al., 2016) dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 7: Qualitative examples of applying DPBridge to semantic segmentation task on CelebAMask-HQ (Lee et al., 2020) dataset. We transform discrete labels into color maps and us our bridge model to predict colors in RGB space instead of categorical labels.
Figure 8: Qualitative examples of applying DPBridge to optical flow prediction task on MPI-Sintel (Butler et al., 2012) dataset. The input image is the overlayed RGB pair, and we transform the flow map into RGB space so that the diffusion bridge process can be constructed accordingly.
Figure 9: Qualitative examples of applying DPBridge to style transfer task on Face-to-Comics (Sxela, 2021) dataset.